In an electric circuit, for driving a current of 3A for 5 minutes, 1200 J of work has to be done. Fine the electromotive force applied on the circuit.

Grade:12

2 Answers

Soumya Ranjan Mohanty
6 years ago
Given: current=I=3A, time=t=5 minutes=5*60=300 seconds, work =W=1200 J
we know amount of charge flowing in the circuit=Q=It=3*300=900 C.
Electromotive force or emf is the work per unit charge in the circuit, or emf=W/Q=1200/900=4/3=1.33 V
So electromotive force applied on circuit is 1.33V.
Thanks,
Soumya Ranjan Mohanty.
Vikas TU
6 years ago
Dear Student,
P=i2R
=>R=P/i2
=>R=1200/9x5
=>R=80/3W
Electromotive force= iR (in Volts)
=80V.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)